# New measurements of the Lyman- forest continuum and effective optical depth with LyCAN and DESI Y1 data

### v1

@@ 5.2 Corrections and Systematic Error
Our measurement of the total optical depth described in §5.1 includes other absorption features in addition to Ly, most notably metal lines and optically thick absorbers. There are also small biases in the continuum predictions, which are illustrated in Figure 12. We correct the total optical depth measurement for both the continuum biases and other absorption features to determine the evolution of the optical depth of the Ly forest. We first use the DESI Y5 mock dataset to calibrate our corrections for the redshift bias in the LyCAN continuum predictions. Figure 14 shows our effective optical depth measurements for the mock dataset based on the LyCAN predicted continua and the mock truth continua. We attribute the difference to the continuum bias, which may be due to the systematic selection of higher luminosity quasars at higher redshifts. The top right panel in Figure 12 shows that LyCAN tends to slightly under-predict the true continuum at lower redshifts and slightly over-predict the continuum at higher redshifts. This effect is small (% in FFE) and is consistent with what is shown in Figure 14 – the under-predictions at lower redshifts lead to less measured absorption, while the over-predictions at high redshifts lead to greater measured absorption than truth. Since the mock quasar samples in Figure 14 are identical, the difference in continuum solely determines the observed discrepancy in measured optical depth.
@@
After correcting our DESI Y1 optical depth measurements for the LyCAN continuum bias, we apply corrections due to absorption from metals and optically thick absorbers. Faucher-Giguère et al. (2008) presented additive corrections due to metal absorption lines per redshift bin using the Schaye et al. (2003) correction scheme. They also calculated a systematic error for this metal correction by examining the deviation between the Schaye et al. (2003) and Kirkman et al. (2005) metal correction schemes. We adopt the metallicity correction from Schaye et al. (2003) and the associated systematic errors from Faucher-Giguère et al. (2008). The systematic error due to the metal correction scheme dominates our total error budget.
@@ 5.3 Comparison with the literature
Previous measurements of the optical depth have either used a relatively small number of high resolution, high SNR spectra or a large number of more moderate resolution spectra with generally lower SNR. The high resolution studies offer the promise of a direct measurement of the continuum between the absorption features, particularly at lower redshifts, while large samples of moderate resolution spectra offer greater statistical power. Faucher-Giguère et al. (2008) performed a direct measurement using 86 high-resolution quasar spectra. In contrast, Becker et al. (2013) used 6,065 moderate resolution SDSS DR7 spectra for a relative measurement of the mean transmitted flux as a fraction of its value at with the assumption that the mean unabsorbed quasar continuum does not evolve with redshift. They used results from Faucher-Giguère et al. (2008) to convert this to an absolute measurement. Kamble et al. (2020) used 40,035 SDSS DR12 quasar spectra to perform a similar relative measurement while also allowing for spectral diversity. They found a higher opacity at compared to Becker et al. (2013). More recently, Liu & Bordoloi (2021) used their feedforward neural network-based continuum predictions on SDSS DR16 quasar spectra to perform an absolute measurement of the effective optical depth. Additionally, Ding et al. (2023) presented a new measurement of the effective Ly forest opacity based on 27,008 SDSS DR14 quasar spectra using a Markov Chain Monte Carlo analysis. The authors found a systematic offset towards lower opacity values compared to Kamble et al. (2020).
@@
Some previous works reported evidence for a decrease in the effective optical depth at . Such a feature could be a signature of He II reionization, as it could signal a rise in the IGM gas temperature at the end of helium reionization (see Theuns et al., 2002; Faucher-Giguère et al., 2008; Becker et al., 2013). Bernardi et al. (2003) found a sudden % reduction in the effective optical depth at this redshift, which was later corroborated by Faucher-Giguère et al. (2008) despite substantial differences in their datasets and analysis methods. However, others have argued from a theoretical perspective that such a sharp feature should not be observed (e.g. Bolton et al., 2009). Recent temperature measurements also suggest that IGM heating due to helium reionization occurred over a more prolonged period, which would not produce a sharp feature in the evolution of the effective optical depth (Becker et al., 2011). Subsequent measurements by Becker et al. (2013) revealed no evidence for a departure from a smooth power-law description of the mean opacity. Later works (e.g. Kamble et al., 2020; Liu & Bordoloi, 2021; Ding et al., 2023) also reported no evidence for this feature.
@@
Our results for the evolution of the mean optical depth show a smooth power-law relationship of the form where and , and no evidence for a dip at . Our approach should be sensitive to this potential feature, as our systematic errors vary smoothly with redshift. We find the closest agreement with measurements based on high-resolution spectra (e.g. Faucher-Giguère et al., 2008), particularly at the lower redshift end. High-resolution spectra are subject to smaller errors in the continuum fitting process, and continuum errors are minimal at lower redshifts where Ly forest absorption is reduced. This agreement indicates that LyCAN can accurately predict the underlying quasar continuum. We also measure a lower optical depth at higher redshifts () compared to previous works, which may indicate a stronger ionizing background at these redshifts. Further, we find good agreement with Liu & Bordoloi (2021), whose measurements were also corrected for metals according to Schaye et al. (2003), over our whole redshift range. This points to the self-consistency of deep learning methods for quasar continuum prediction. Our measurements are shown alongside others in Figure 15. The error bars for the Becker et al. (2013) and Liu & Bordoloi (2021) measurements are likely underestimated as they do not include the systematic error in the correction for metal line absorption.
@@ 6 Summary
We used the LyCAN continuum predictions on DESI Y1 quasars to perform an absolute measurement of the evolution of the effective optical depth with the largest sample to date. Our results show close agreement with previous measurements based on high-resolution quasar spectra at lower redshifts (e.g. Faucher-Giguère et al., 2008) where those methods are most able to measure the quasar continuum between the absorption features. This demonstrates that LyCAN can accurately predict the unabsorbed continuum. We find that the mean optical depth varies smoothly with redshift, and do not find evidence for a feature at that could be associated with He II reionization. Due to our large sample size of 83,635 quasars and LyCAN’s ability to accurately predict the underlying continuum, we expect that this is the most accurate measurement of the effective optical depth evolution to date.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
@@
The data points for each figure are available at https://doi.org/10.5281/zenodo.11106582 (catalog doi:10.5281/zenodo.11106582).

### v2

@@ 5.2 Corrections and Systematic Error
Our measurement of the total optical depth described in §5.1 includes other absorption features in addition to Ly, most notably metal lines and optically thick absorbers. There are also small biases in the continuum predictions, which are illustrated in Figure 12. We correct the total optical depth measurement for both the continuum biases and other absorption features to determine the evolution of the optical depth of the Ly forest. We first use the DESI Y5 mock dataset to calibrate our corrections for the redshift bias in the LyCAN continuum predictions. Figure 14 shows our effective optical depth measurements for the mock dataset based on the LyCAN predicted continua and the mock truth continua. We attribute the difference to the continuum bias, which may be due to the systematic selection of higher luminosity quasars at higher redshifts. The top right panel in Figure 12 shows that LyCAN tends to slightly under-predict the true continuum at lower redshifts and slightly over-predict the continuum at higher redshifts. This effect is small (% in FFE) and is consistent with what is shown in Figure 14 â the under-predictions at lower redshifts lead to less measured absorption, while the over-predictions at high redshifts lead to greater measured absorption than truth. Since the mock quasar samples in Figure 14 are identical, the difference in continuum solely determines the observed discrepancy in measured optical depth.
@@
After correcting our DESI Y1 optical depth measurements for the LyCAN continuum bias, we apply corrections due to absorption from metals and optically thick absorbers. Faucher-GiguÚre et al. (2008) presented additive corrections due to metal absorption lines per redshift bin using the Schaye et al. (2003) correction scheme. They also calculated a systematic error for this metal correction by examining the deviation between the Schaye et al. (2003) and Kirkman et al. (2005) metal correction schemes. We adopt the metallicity correction from Schaye et al. (2003) and the associated systematic errors from Faucher-GiguÚre et al. (2008). The systematic error due to the metal correction scheme dominates our total error budget.
@@ 5.3 Comparison with the literature
Previous measurements of the optical depth have either used a relatively small number of high-resolution, high SNR spectra or a large number of more moderate resolution spectra with generally lower SNR. The high-resolution studies offer the promise of a direct measurement of the continuum between the absorption features, particularly at lower redshifts, while large samples of moderate resolution spectra offer greater statistical power. Faucher-GiguÚre et al. (2008) performed a direct measurement using 86 high-resolution quasar spectra. In contrast, Becker et al. (2013) used 6,065 moderate resolution SDSS DR7 spectra for a relative measurement of the mean transmitted flux as a fraction of its value at with the assumption that the mean unabsorbed quasar continuum does not evolve with redshift. They used results from Faucher-GiguÚre et al. (2008) to convert this to an absolute measurement. Kamble et al. (2020) used 40,035 SDSS DR12 quasar spectra to perform a similar relative measurement while also allowing for spectral diversity. They found a higher opacity at compared to Becker et al. (2013). More recently, Liu & Bordoloi (2021) used their feedforward neural network-based continuum predictions on SDSS DR16 quasar spectra to perform an absolute measurement of the effective optical depth. Gaikwad et al. (2021) used a sample of 103 high-resolution, high SNR quasar spectra from kodiaq dr2 to directly measure the evolution of the mean flux, and found results in agreement with Faucher-GiguÚre et al. (2008). Additionally, Ding et al. (2023) presented a new measurement of the effective Ly forest opacity based on 27,008 SDSS DR14 quasar spectra using a Markov Chain Monte Carlo analysis. The authors found a systematic offset towards lower opacity values compared to Kamble et al. (2020).
@@
Some previous works reported evidence for a decrease in the effective optical depth at . Such a feature could be a signature of He II reionization, as it could signal a rise in the IGM gas temperature at the end of helium reionization (see e.g. Theuns et al., 2002; Faucher-GiguÚre et al., 2008; Becker et al., 2013). Bernardi et al. (2003) found a sudden % reduction in the effective optical depth at this redshift, which was later corroborated by Faucher-GiguÚre et al. (2008) despite substantial differences in their datasets and analysis methods. However, others have argued from a theoretical perspective that such a sharp feature should not be observed (e.g. Bolton et al., 2009). Some temperature evolution measurements (e.g. Becker et al., 2011; Walther et al., 2019) also suggest that IGM heating due to He II reionization occurred over a more prolonged period, which would not produce a sharp feature in the evolution of the effective optical depth. However, more recent temperature measurements from Gaikwad et al. (2021) provide evidence for a more rapid He II reionization. Additional measurements of the mean transmitted flux by Becker et al. (2013) revealed no evidence for a departure from a smooth power-law description. Later works (e.g. Kamble et al., 2020; Liu & Bordoloi, 2021; Ding et al., 2023) also reported no evidence for this feature.
@@
Our results for the evolution of the mean optical depth show a smooth power-law relationship of the form where and , and no evidence for a dip at . Our approach should be sensitive to this potential feature, as our systematic errors vary smoothly with redshift. We find the closest agreement with measurements based on high-resolution spectra (e.g. Faucher-GiguÚre et al., 2008; Gaikwad et al., 2021), particularly at the lower redshift end. High-resolution spectra are subject to smaller errors in the continuum fitting process, and continuum errors are minimal at lower redshifts where Ly forest absorption is reduced. This agreement indicates that LyCAN can accurately predict the underlying quasar continuum. We also measure a lower optical depth at higher redshifts () compared to previous works, which may indicate a stronger ionizing background at these redshifts. Further, we find good agreement with Liu & Bordoloi (2021), whose measurements were also corrected for metals according to Schaye et al. (2003), over our whole redshift range. This points to the self-consistency of deep learning methods for quasar continuum prediction. Our measurements are shown alongside others in Figure 15. The error bars for the Becker et al. (2013) and Liu & Bordoloi (2021) measurements are likely underestimated as they do not include the systematic error in the correction for metal line absorption.
@@ 6 Summary
We used the LyCAN continuum predictions on DESI Y1 quasars to perform an absolute measurement of the evolution of the effective optical depth with the largest sample to date. Our results show close agreement with previous measurements based on high-resolution quasar spectra at lower redshifts (e.g. Faucher-GiguĂšre et al., 2008; Gaikwad et al., 2021) where those methods are most able to measure the quasar continuum between the absorption features. This demonstrates that LyCAN can accurately predict the unabsorbed continuum. We find that the mean optical depth varies smoothly with redshift, and do not find evidence for a feature at that could be associated with He II reionization. Due to our large sample size of 83,635 quasars and LyCANâs ability to accurately predict the underlying continuum, we expect that this is the most accurate measurement of the effective optical depth evolution to date.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DEâAC02â05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSFâs National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
The authors are honored to be permitted to conduct scientific research on Iolkam Duâag (Kitt Peak), a mountain with particular significance to the Tohono Oâodham Nation.
@@
The data points for each figure are available at https://doi.org/10.5281/zenodo.11106581 (catalog doi:10.5281/zenodo.11106581). The COS observations used for generation of synthetic spectra can be accessed via https://doi.org/10.17909/ae35-e630 (catalog doi: 10.17909/ae35-e630).